1. Phi Upsilon Nu, a student social organization, has two different locations under consideration for constructing a new chapter house. PhUN's president, a POM student, estimates that due to differing land costs, utility rates, etc., both fixed and variable costs would be different for each of the proposed sites, as follows: ANNUAL OPERATING COSTS LOCATION FIXED VARIABLE Alpha Ave. $5,000 $200 per person Beta Blvd. $8,000 $150 per person What would be total annual costs for either location at the point of indifference? Question 20 options: $13,000 $13,350 $9,000 $17,000 $19,200

2.

The production planner for Fine Coffees, Inc. produces two coffee blends: American (A) and British (B). Two of his resources are constrained: Columbia beans, of which he can get at most 300 pounds (4,800 ounces) per week; and Dominican beans, of which he can get at most 200 pounds (3,200 ounces) per week. Each pound of American blend coffee requires 12 ounces of Colombian beans and 4 ounces of Dominican beans, while a pound of British blend coffee uses 8 ounces of each type of bean. Profits for the American blend are $2.00 per pound, and profits for the British blend are $1.00 per pound. Which of the following is not a feasible production combination?

Question 15 options:

0 A & 0 B

0 A & 400 B

200 A & 300 B

400 A & 0 B

400 A & 400 B

3.

Which of the choices below constitutes a simultaneous solution to these equations (1) 3x + 4y = 10 and, (2) 5x + 4y = 14?

Question 14 options:

x = 2, y = .5

x = 4, y = -.5

x = 2, y = 1

x = y

y = 2x

4.

The Production Line is in operation for 7.0 hours per day. Desired output is 210 units per day. To achieve this level of output the Cycle time in minutes is

5. A production line is to be designed for a job with three tasks. The task times are 0.5 minutes, 0.8 minutes, and 1.2 minutes. The minimum cycle time in minutes is

6.

A production line is to be designed for a job with three tasks. The task times are 0.7 minutes, 2.1 minutes, and 0.3 minutes. The maximum cycle time in minutes is
Answer #1

Q1 . Point of indifference = (Difference in fixed costs ) / (Difference in variable costs) = (8000-5000) / 50 = 60 units

Cost at Point of indifference = 8000 + 60*150 = $17000 or 5000+200*60 = $17000

Q2 400 A & 400 B is not a feasible solution

For 400 each, the Colombian beans required = 400*12+400*8 = 8000 ounces. Our constraint is 3200 ounces

Note - Solved Q1 & Q2 as per Q&A policy. Please post unrelated questions separately.